# MULAN-WC: Multi-Robot Localization Uncertainty-aware Active NeRF with Wireless Coordination

### v1

@@ III-B Uncertainty-aware NeRF training
As described in our problem formulation, if we have two sets of camera poses $T_{\alpha}$ and $T_{\beta}$ in their own local frames, we want to find the relative transformation between them to align the poses in the same frame of reference. For each wireless measurement between measuring poses $T_{k}^{\alpha}$ and $T_{p}^{\beta}$, we can quantify the uncertainty of the AoA and ranging estimates using the methods described in the previous section.
Let $\sigma_{k,p}$ and $\kappa_{k,p}$ denote the variance of the ranging measurement and the AoA, respectively, for the wireless measurement between poses $T_{k}^{\alpha}$ and $T_{p}^{\beta}$. We propose incorporating these uncertainty measures into the NeRF training process by modifying the standard pixel loss function $\mathcal{L}$ given in Eq 1 by re-scaling the loss for each training sample. For brevity of the annotation in this section, we omit robot indices $\alpha$, $\beta$, and all local time frames p and k from now on. We apply uncertainty propagation to compute the error ellipse of uncertain measurements. Since the AoA measurements and ranging measurements are taken using different sensing modalities, they can be treated as independent measurements with the error ellipse’s axes aligning with x-y axes. The semi-major and semi-minor axes $a$ and $b$ of the error ellipse can be derived as:
@@
where $t$ is the ranging measurement, and $\theta$ is the AoA estimate. Then the scale factor with confidence interval $CI$ is given by
@@
[…]
@@ III-C Active Best-View finding with position uncertainty
In a multi-robot NeRF system, actively selecting the best views for each robot to capture can significantly improve the efficiency and quality of the 3D reconstruction. However, the uncertainty in robot poses obtained through wireless coordination can impact the effectiveness of the view selection process. When a robot attempts to find the best next view location by proposing and evaluating potential new positions, the uncertainty in its current pose can lead to inaccurate assessments of the information gain at these novel view locations.
To address this challenge, we propose an active view finding approach that incorporates the position uncertainty of the robots to guide the selection of the most informative views for NeRF training. Building upon the approach proposed in [1] for evaluating the potential information gain from novel views by quantifying the reduction of variance in rendering, we extend this method to account for the uncertainty in the robot’s current position and its propagation to the novel view locations being evaluated. By considering the localization uncertainty during the novel view selection process, we can more accurately determine the most informative next views for each robot, even in the presence of pose uncertainty arising from wireless coordination.
We adopt the assumption that the radiance color of any location along the ray $r(t)$ can be parameterized by a Gaussian distribution with mean $\bar{c}(r(t))$ and variance $\bar{\beta}(r(t))$. To incorporate this uncertainty, we model the origin $\boldsymbol{o}$ of each ray $r$ following a Gaussian distribution with 0 mean and the variance $\sigma$, representing the localization uncertainty.
@@
Assuming a NeRF model $\mathcal{M}$ has been trained on an initial collection of data $D$, the prior distribution $P(c(r(t_{k}))|D)$ of the color $c$ at location $r(t)$ follows a Gaussian distribution $\mathcal{N}\sim(\bar{c}(r(t)),\bar{\beta}^{2}(r(t))$. The accumulated color from a new ray r passing through can also be modeled as a Gaussian distribution:
@@
$p(C(r)|c(r(t))$ $=\int p(C(r)|c(r(t_{k}))*p(\boldsymbol{o})d\boldsymbol{o}$ (13)
$p(C(r)|c(r(t_{k})$ $\sim\mathcal{N}(\sum_{i=1}^{N}\alpha_{i}\bar{c}(r(t),\sum_{i=1}^{N}\alpha_{i}*\sigma^{2}+\bar{\beta}^{2}(r(t_{k}))$ (14)
@@
$\propto P(C(r)|c(r(t_{k}))*P(c(r(t_{k}))|D)*P(\boldsymbol{o})$
$*\displaystyle\propto exp(-\frac{1}{2}*(c(r(t_{k}))-(\frac{\omega C(r)}{\alpha_{k}}+(1-\omega))*\bar{c}(r(t_{k})))*$
$(\frac{\alpha_{k}^{2}}{\alpha_{k}^{2}\sigma^{2}+\bar{\beta}^{2}(r)}+\frac{1}{\bar{\beta}^{2}(r(t_{k}))})^{-1})$
@@
Then the variance of the posterior distribution can be extracted as
@@
As the localization variance increases, the uncertainty of the radiance field also increases accordingly. To select the best view, the metric will prefer the novel view with lower localization uncertainty and lower Since we only need to consider the variance reduction given multiple rays from a sampled novel-view position, we can then command the robot to move to the location with highest variance reduction using Equation 16.
@@ IV Results
In this section, we present a comprehensive evaluation of our methodology through both synthetic datasets collected in synthetic environments, and real-world datasets collected on our hardware robots. Our findings validate the effectiveness of our algorithm in integrating perspectives from multiple robots within an active acquisition framework, showcasing significant improvements in data capture and processing. We implemented our algorithm based on a Pytorch implementation of the SDF and NeRF part described in Instant-NGP [11] with CUDA-accelerated ray marching. We modified the loss function to implement the localization uncertainty-aware loss described in Equation 1. Furthermore, for the active image collection, we re-implemented the rendering variance reduction described in [17] incorporated with the localization uncertainty as described in Eq 16. We use a desktop with NVIDIA RTX 6000 for all evaluations.
@@ IV-A Wireless variance
First, we present the performance benchmark with our proposed wireless variance metric as formulated in Section III-A. With a simulated trajectory, the metric is tested over $16$ random trials with various amounts of injected Gaussian channel noise as defined in Equation 5, ranging from $0.01$ to $3$ radians phase noise standard deviation. As shown in Figure 2, the AoA error from the ground truth scales up quickly and becomes unstable as our proposed AoA variance metric increases. To further demonstrate the relationship between the variance of the AoA Error and our proposed profile variance, Figure 3 clearly shows that the variance of the AoA error will exponentially increase as our metric increases. This result indicates that our AoA uncertainty quantification is well aligned as an indicator of the variance of AoA measurements, which can be further used to quantify the uncertainty of the camera position.
Figure 2: The Absolute AoA error from ground truth plotted against our AoA uncertainty metric. We see that nonzero AoA error grows as our AoA uncertainty metric grows. This indicates that our metric successfully captures true error in measured AoA.
@@ IV-B Simulation Experiment
Our algorithm localization uncertainty-aware framework described in Sec III-B is first assessed using a synthetic dataset lego released with the original NeRF work [19]. The dataset is commonly used for evaluating NeRF frameworks. The dataset is partitioned into two subsets to simulate the data acquired from two robots. This approach allows us to mimic the real-world scenario of capturing images from different angles and positions, thereby testing the robustness and adaptability of our algorithm in synthesizing and analyzing data from varied viewpoints. There are total 3 setups are evaluated:
@@
    [Oracle]: Camera poses from the dataset in the global frame (known as extrinsic between cameras) and images from the dataset
@@
    [Normalized camera poses with AOA and ranging simulated]]: Camera poses from the dataset but normalized by the first pose in each partition. Then the AOA and ranging are simulated with noise whose standard deviation are 0.05 meters and 5 degrees respectively.
@@
    [Normalized camera poses with AOA and ranging simulated with variance as supervision]: The same setup in B but the training loss is incorporated with the noise variance.
@@
TABLE I: Performance comparison between different setups where larger PSNR values are better, and smaller LPIPS values indicate better quality. The comparison demonstrate that applying the uncertainty-aware loss can effectively improve the quality of the model
@@ IV-C Hardware Experiment
For the real-world application, we deployed our algorithm on two of customized Locobot PX100 robots. These robots were equipped with Oak-D Pro Cameras, operating at 1080p 20Hz, along with DWM1001 UWB modules, 5dBi Antennas, and Intel NUC 10 computers for onboard processing. The experimental setup involves a drone object centrally relative to the two robots, which are programmed to navigate curved paths around the object to complete data capture.
Both robots utilize onboard Visual Inertial Odometry (VIO) to estimate local camera displacement within their respective frames. At the onset of the experiment, Angle of Arrival (AoA) and ranging measurements are taken to establish an initial estimate of the relative positioning between the robots. Subsequently, the covariance of the VIO data was monitored to identify optimal intervals for refreshing wireless data collection. In the meantime, the testbed is equipped with the Optitrack motion capture system providing the ground truth camera poses for each robot.
@@
All five setups are evaluated using standard metrics for NeRF: Peak Signal-to-Noise Ratio (PSNR) and Learned Perceptual Image Patch Similarity (LPIPS)[22]. Each metric is evaluated from samples in the test set ground truth images, along with camera poses. For each setup, there is a total of 100 images with camera poses that are collected continuously from each robot while robots are moving around the drone subject. drone-1 an drone-2 are different images in the testing dataset.
@@
TABLE II: Performance comparison between different setups where larger PSNR values are better, and smaller LPIPS values indicate better quality. The comparison between setup A and setup B demonstrates that applying wireless coordination can effectively achieve close performance to having a global coordinate system. Results from setup C show the realistic performance of our system using fully onboard VIO for local positioning, which is degraded but still relatively robust. Setup D shows the scaling with localization uncertainty quantification can improve the quality almost to the best-case scenario in setup B.
As shown in Table II which provides our quantitative results, setup A shows the best performance we can achieve in a two-robot team since it is based on the ground truth camera poses provided by the motion capture system. In setup B, the poses are normalized by the starting pose of each robot captured in the motion capture system. Then wireless coordination is incorporated to provide inter-robot camera extrinsic. Setup C shows the realistic setup, which applies the wireless localization between robots to the local VIO poses and is effectively close to the result in setup B. Setup D shows that our framework can achieve better results than C by using the variance-aware loss function defined in Equation 10 with the corresponding localization variance proposed in Equation 16. The benchmark comparison setup E fails to produce a cohesive 3D rendering, due to the discrepancy of the relative camera pose estimation using COLMAP [21]. This is mainly due to the drastic translation change in camera view from different robots.
Figure 4: PNSR improvement over iterations with different setups
In many robotics applications, there are possibilities that we need to have a quickly-converging view of the environment before the model training fully converges. In our experiment, we also validate that our methods not only deliver better rendering but also achieve faster PSNR improvement as shown in Fig 4.
@@ IV-D Active Image Capturing
Following the initial phase of data gathering, a waiting period is observed until the Neural Radiance Field (NeRF) model’s loss stabilized. The robots then execute a series of maneuvers, sampling eight different directions at 0.5-meter intervals. Our evaluation focused on minimizing variance of the rendering posterior, employing Equation 16 to identify positions yielding the most significant reduction in variance. The application of our algorithm in a hardware setting demonstrates its practical feasibility. Moreover, it underscores the potential of our method to optimize the data capture process through strategic robot positioning and movement.
After selecting the location with the highest variance reduction using our proposed method, the robot is commanded to the new location and observes the environment again. We then let the model train until the loss stabilizes and repeat the process four times to evaluate the efficacy of our method. For comparison, we also randomly selected accessible locations around the robots and controlled the robots to move to those locations. The evaluation-maneuver-training loop was conducted on both our policy and the random policy and the results are reported in Table III These results demonstrate that our approach provides a principle metric that can improve the quality of the rendering consistently.
@@ V Conclusion
This work presents MULAN-WC, a multi-robot 3D reconstruction method that uses wireless signal-based coordination between robots. This work presents i) a framework for multi-robot NeRF that uses SAR-based wireless relative position measurements to stitch together views of the environment from multiple robots, ii) uncertainty-based weighting of samples in the NeRF training as a supervision technique, where samples with greater wireless measurement noise are weighted less, leading to better accuracy of the combined rendering, and iii) collaborative active next-image acquisition, where novel-view location sampling incorporates wireless pose uncertainty, and is used to direct robots to better sampling locations that reduce variance during NeRF training. We demonstrate the performance of the multi-robot framework in hardware, where our results show good quality of rendering according to the standard NeRF error metrics of PSNR and LPIPS, and an consistent improvement when we additionally use the uncertainty of the AoA measurement as supervision in the NeRF training. Lastly, we show that AoA measurements can be used to select next-best-view based on regions of better position accuracy and that this results in incremental rendering quality improvement.
@@ Acknowledgments
The authors gratefully acknowledge partial funding through NSF grant ##\##CNS-2114733 and the Sloan award ##\##FG-2020-13998.

### v2

@@ III-B Uncertainty-aware NeRF Training
As described in our problem formulation, if we have two sets of camera poses $T_{\alpha}$ and $T_{\beta}$ in their own local frames, we want to find the relative transformation between them to align the poses in the same frame of reference. For each wireless measurement between poses $T_{k}^{\alpha}$ and $T_{p}^{\beta}$, we can quantify the uncertainties $\sigma_{k,p}$ and $\kappa_{k,p}$ respectively of the ranging and AoA estimates using the methods described in the previous section. We propose incorporating these uncertainty measures into the NeRF training process by modifying the standard pixel loss function $\mathcal{L}$ given in Eq 1 by re-scaling the loss for each training sample. For brevity of the annotation in this section, we omit robot indices $\alpha$, $\beta$, and all local time frames p and k from now on. We apply uncertainty propagation to compute the error ellipse of uncertain measurements. Since the ranging and AoA measurements are taken using different sensing modalities, they can be treated as independent measurements with the error ellipse’s axes aligning with x-y axes. The semi-major and semi-minor axes $a$ and $b$ of the error ellipse can be derived as:
@@
where $t$ is the ranging measurement and $\theta$ is the AoA estimate. Then the scale factor with confidence interval $CI$ is given by
@@
$\mathcal{L}_{uncertainty}=\textit{SIGMOID}(\gamma)\mathcal{L}$ (10)
@@ III-C Active Best-View Finding with Position Uncertainty
In a multi-robot NeRF system, actively selecting the best views for each robot to capture can significantly improve the efficiency and quality of the 3D reconstruction. However, the uncertainty in robot poses obtained through wireless coordination can impact the effectiveness of the view selection process. When a robot attempts to find the best next view location by proposing and evaluating potential new positions, the uncertainty in its current pose can lead to inaccurate assessments of the information gain at novel view locations.
To address this challenge, we propose an active view finding approach that incorporates the positional uncertainty of the robots to guide the selection of the most informative views for NeRF training. Building upon the approach proposed in [19] for evaluating the potential information gain from novel views by quantifying the reduction of variance in rendering, we extend this method to account for the uncertainty in the robot’s current position and its propagation to the novel view locations being evaluated. By considering localization uncertainty during the novel view selection process, we can more accurately determine the most informative next views for each robot, even in the presence of pose uncertainty arising from wireless coordination.
We adopt the assumption that the radiance color of any location along the ray $r(t)$ can be parameterized by a Gaussian distribution with mean $\bar{c}(r(t))$ and variance $\bar{\beta}(r(t))$. To incorporate this uncertainty, we model the origin $\boldsymbol{o}$ of each ray $r$ following a Gaussian distribution with 00 mean and the variance $\sigma$, representing the localization uncertainty.
@@
Assuming a NeRF model $\mathcal{M}$ has been trained on an initial collection of data $D$, the prior distribution $P(c(r(t_{k}))|D)$ of the color $c$ at location $r(t)$ follows a Gaussian distribution $\mathcal{N}\sim(\bar{c}(r(t)),\bar{\beta}^{2}(r(t)))$. The accumulated color from a new ray $r$ passing through can also be modeled as a Gaussian distribution:
@@
[…]
$p(C(r)|c(r(t_{k})))$ $\sim\mathcal{N}(\sum_{i=1}^{N}\alpha_{i}\bar{c}(r(t)),\sum_{i=1}^{N}\alpha_{i}*\sigma^{2}+\bar{\beta}^{2}(r(t_{k})))$ (14)
@@
$\propto P(C(r)|c(r(t_{k})))*P(c(r(t_{k}))|D)*P(\boldsymbol{o})$
[…]
$\left.\left(\frac{\alpha_{k}^{2}}{\alpha_{k}^{2}\sigma^{2}+\bar{\beta}^{2}(r)}+\frac{1}{\bar{\beta}^{2}(r(t_{k}))}\right)^{-1}\right)$
@@
Then extract the variance of the posterior distribution:
@@
As the localization variance increases, the uncertainty of the radiance field also increases accordingly. To select the best view, the metric will prefer the novel view with lower localization uncertainty. Since we only need to consider the variance reduction given multiple rays from a sampled novel-view position, we can then command the robot to move to the location with highest variance reduction using Equation 16.
@@ IV Results
In this section, we present a comprehensive evaluation of our methodology through both synthetic datasets collected in synthetic environments and real-world datasets collected on our hardware robots. Our findings validate the effectiveness of our algorithm in integrating perspectives from multiple robots within an active acquisition framework, showcasing significant improvements in data capture and processing. We implemented our algorithm based on a Pytorch implementation of the SDF and NeRF part described in Instant-NGP [11] with CUDA-accelerated ray marching. We modified the loss function to implement the localization uncertainty-aware loss described in Eq 1. Further, for the active image collection, we re-implemented the rendering variance reduction described in [19], incorporating the localization uncertainty from Eq 16. We use a desktop with NVIDIA RTX 6000 for all evaluations.
@@ IV-A Wireless Variance
First, we present the performance benchmark with our proposed wireless variance metric as formulated in Section III-A. With a simulated trajectory, the metric is tested over $16$ random trials with various amounts of injected Gaussian channel phase noise as defined in Equation 5, ranging from $0.01$ to $3$ radians standard deviation. As shown in Figure 2, the AoA error from the ground truth scales quickly and becomes unstable as our proposed AoA variance metric increases. To further demonstrate the relationship between the variance of the AoA error and our proposed profile variance, Figure 3 clearly shows that the variance of the AoA error will increase superlinearly as our metric increases. This result indicates that our AoA uncertainty quantification is well aligned as an indicator of the variance of AoA measurements, which can be further used to quantify the uncertainty of the camera position.
Figure 2: Absolute AoA error from ground truth plotted against our AoA uncertainty metric. We see that nonzero AoA error grows as our AoA uncertainty metric grows, indicating that our metric successfully captures true error in measured AoA.
@@ IV-B Simulation Experiment
Our localization uncertainty-aware framework described in Sec III-B is first assessed using a synthetic dataset lego released with the original NeRF work [21] that is commonly used for evaluating NeRF frameworks. The dataset is partitioned into two subsets to simulate data acquired from two robots, allowing us to mimic the real-world scenario of capturing images from different angles and positions, thereby testing the robustness and adaptability of our algorithm in synthesizing and analyzing data from varied viewpoints. Three setups are evaluated with results in Table I:
@@
    [Oracle]: Camera poses from the dataset in the global frame (known as extrinsic between cameras) and images from the dataset.
@@
    [Normalized camera poses with AOA and ranging simulated]: Camera poses from the dataset but normalized by the first pose in each partition. Then the AOA and ranging are simulated with noise whose standard deviation are 0.05 meters and 5 degrees respectively.
@@
    [Normalized camera poses with AOA and ranging simulated with variance as supervision]: The same setup as B but the training loss is incorporated with the localization variance.
As predicted, training using localization variance produces better performance closer to using ground truth poses.
@@
TABLE I: Performance comparison between different setups where larger PSNR values are better, and smaller LPIPS values indicate better quality, illustrating that applying uncertainty-aware loss can effectively improve the quality of the model.
@@ IV-C Hardware Experiment
For a real-world application, we deployed our algorithm on two customized Locobot PX100 robots. These robots were equipped with Oak-D Pro cameras, operating at 1080p 20Hz, along with DWM1001 UWB modules, 5dBi Antennas, and Intel NUC 10 computers for onboard processing. The experimental setup places a drone as a test object central relative to the two robots, which are programmed to navigate curved paths around the object to complete data capture. The wireless AOA measurements are computed by deploying [17] which only requires very small communication bandwidth at around 5 kB/s.
Both robots utilize onboard Visual Inertial Odometry (VIO) to estimate local camera displacement within their respective frames. At the onset of the experiment, Angle of Arrival (AoA) and ranging measurements are taken to establish an initial estimate of the relative positioning between the robots. Subsequently, the covariance of the VIO data was monitored to identify optimal intervals for refreshing wireless data collection. In the meantime, the testbed is equipped with the Optitrack motion capture (mocap) system providing the ground truth camera poses for each robot.
@@
All five setups are evaluated using standard metrics for NeRF: Peak Signal-to-Noise Ratio (PSNR) and Learned Perceptual Image Patch Similarity (LPIPS) [23]. Each metric is evaluated from samples in the test set ground truth images, along with camera poses. For each setup, there are a total of 100 images with camera poses that are collected continuously from each robot while robots are moving around the drone subject. drone-1 an drone-2 are different images in the testing dataset.
@@
TABLE II: Performance comparison between different setups where larger PSNR values are better, and smaller LPIPS values indicate better quality. The comparison between setup A and setup B demonstrates that applying wireless coordination can effectively achieve close performance to having a global coordinate system. Results from setup C show the realistic performance of our system using fully onboard VIO for local positioning, which is degraded but still relatively robust. Setup D shows the scaling with localization uncertainty quantification can improve the quality almost to the best-case scenario in setup B. Setup E fails to produce a coherent 3D rendering.
As shown in Table II which provides our quantitative results, setup A shows the best performance we can achieve in a two-robot team since it is based on the ground truth camera poses provided by the motion capture system. In setup B, the poses are normalized by the starting pose of each robot captured in the motion capture system. Then wireless coordination is incorporated to provide inter-robot camera extrinsic. Setup C shows the realistic setup, which applies the wireless localization between robots to the local VIO poses and is effectively close to the result in setup B. Setup D shows that our framework can achieve better results than C by using the variance-aware loss function defined in Equation 10 with the corresponding localization variance proposed in Equation 16. The benchmark comparison setup E fails to produce a cohesive 3D rendering due to discrepancies in the relative camera pose estimation using COLMAP [22], which is commonly used for estimating the relative camera pose given two frames from different views. This is mainly due to the drastic translation change in camera view from different robots. This results also suggests that COLMAP won’t be a proper solution for a multi-robot setup.
Figure 4: PSNR improvement over epochs, with different setups.
Many robotics applications require a quickly-converging view of the environment before the model training fully converges. In our experiment, we also validate that our methods not only deliver better rendering but also achieve faster PSNR improvement as shown in Fig 4.
@@ IV-D Active Image Capturing
For evaluating active best next view, following the initial phase of data gathering, a waiting period was observed until the Neural Radiance Field (NeRF) model’s loss stabilized. The robots then execute a series of maneuvers, sampling eight different directions at 0.5-meter intervals. Our evaluation focused on minimizing the variance of the rendering posterior, employing Equation 16 to identify positions yielding the most significant reduction in variance. The application of our algorithm in a hardware setting demonstrates its practical feasibility. Moreover, it underscores the potential of our method to optimize the data capture process through strategic robot positioning and movement.
After selecting the location with the highest variance reduction using our proposed method, the robot is commanded to the new location and observes the environment again. We then let the model train until the loss stabilizes and repeat the process four times to evaluate the efficacy of our method. For comparison, we also randomly selected accessible locations around the robots and controlled the robots to move to those locations. The evaluation-maneuver-training loop was conducted on both our policy and the random policy and the results are reported in Table III. These results demonstrate that our approach provides a principle metric that can improve the quality of the rendering consistently.
@@ V Conclusion
This work presents MULAN-WC, a multi-robot 3D reconstruction method that uses wireless signal-based coordination between robots. This work presents i) a framework for multi-robot NeRF that uses SAR-based wireless relative position measurements to stitch together views of the environment from multiple robots, ii) uncertainty-based weighting of samples in the NeRF training as a supervision technique, where samples with greater wireless measurement noise are weighted less, leading to better accuracy of the combined rendering, and iii) collaborative active next-image acquisition, where novel-view location sampling incorporates wireless pose uncertainty, and is used to direct robots to better sampling locations that reduce variance during NeRF training. We demonstrate the performance of the multi-robot framework in hardware, where our results show good quality of rendering according to the standard NeRF error metrics of PSNR and LPIPS, and consistent improvement when we additionally use the uncertainty of the AoA measurement as supervision in the NeRF training. Lastly, we show that AoA measurements can be used to select the best-next-view based on regions of better position accuracy and that this results in incremental rendering quality improvement.
@@ Acknowledgments
The authors gratefully acknowledge partial funding through NSF grant ##\##CNS-2114733, Amazon ARA, and the Sloan award ##\##FG-2020-13998.